OpenStudy (anonymous):

What is the square root of pie^2l

OpenStudy (anonymous):

pi

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

It pie to the power of two times l

OpenStudy (anonymous):

square rooting is the removal of squared so it will become Pie as answer

OpenStudy (anonymous):

No

OpenStudy (anonymous):

What happens to the l?

OpenStudy (anonymous):

Pi^21 = 27551631842.9 Square root of 27551631842.9 = 165986.842379

OpenStudy (anonymous):

Hold on, is it Pi^21?

OpenStudy (anonymous):

No it's pi^2L

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

square root of x=x^1/2 square root of pi^2l = pi^(2l/2)=pi^l

OpenStudy (anonymous):

Ohh ok

geerky42 (geerky42):

\[\pi^2 l \Longrightarrow\sqrt{\pi^2l} = \pi\sqrt{l}\] \[\pi^{2l} \Longrightarrow\sqrt{\pi^{2l}} = \sqrt{(\pi^l)^2}=\pi^l\]